OpenStudy (anonymous):

Which property is represented in this equation? 4 · (4 · 2) = (4 · 4) · 2 Distributive Property Commutative Property(i think its this one) Associative Property Identity Property

OpenStudy (anonymous):

Commutative. Yes.

OpenStudy (anonymous):

thank you

myininaya (myininaya):

associative

OpenStudy (anonymous):

yea i got it wrong already haha but thank you for answering was still bugging me

myininaya (myininaya):

commutative is a*b=b*a associative is (a*b)*c=a*(b*c)